# rowhighlight linewidth=1pt, linecolor=black, innertopmargin=4pt, innerbottommargin=4pt, innerleftmargin=4pt, innerrightmargin=4pt, backgroundcolor=lightgray!15

### v1

@@ 4.3 Absorption
The primary challenge in interpretability is feature absorption, where one learned feature becomes a strict subset of another (e.g. a “Lion” feature is entirely subsumed by a “starts with L” feature) and consequently fails to activate on instances that satisfy the broader concept but not its superset representation (Chanin et al., 2024).
Figure 4 reports three absorption metrics across sparsity levels $\ell_{0}\in\{16,32,64,128,256\}$: (1) the mean absorption fraction, measuring the proportion of features that are partially absorbed; (2) the mean full-absorption score, quantifying complete subsumption events; and (3) the mean number of feature splits, indicating how often a single conceptual feature fragments into multiple activations. Across all $\ell_{0}$, KronSAE variants consistently reduce both the absorption fraction and the full-absorption score relative to the TopK SAE baseline, while maintaining a similar rate of feature splits. We attribute KronSAE’s improved disentanglement to two complementary design choices:
@@
    Smooth $\operatorname{mAND}$ activation. By emitting nonzero output only when both pre-latents are positive, we introduce a differentiable AND gate that prevents a broadly polysemantic primitive from entirely subsuming a more specific one. Consequently, composite post-latents fire mainly in the intersection of their constituent concepts, encouraging each pre-latent to specialize on a single semantic mode rather than inherit its “parent” activation region.
@@
    Head-wise Cartesian decomposition. Dividing the latent space into $h$ independent subspaces (each with its own $m\times n$ grid of primitive interactions) ensures that specialized concepts (such as “elephant”) are confined to a single head and cannot fully absorb more general concepts (such as “starts with E”) in another.
Figure 3: (a) EV versus sparsity level $L_{0}$ for KronSAE and TopK SAE on the 14th layer of Qwen-2.5-1.5B and the 12th layer of Gemma-2-2B under iso-FLOPs constraints. (b) EV across layers of Qwen-2.5-1.5B, demonstrating that KronSAE matches TopK performance regardless of depth.
Together, these mechanisms produce more monosemantic features, simplifying downstream causal interventions and targeted probing. Notably, the mean number of feature splits remains comparable to the TopK baseline, as Cartesian decomposition alone does not inherently alter the fragmentation of individual primitives (see Section 5.2).
Figure 4: Feature absorption metrics on Qwen-2.5-1.5B. KronSAE configurations (various $m,n$) exhibit lower mean absorption fractions and full-absorption scores across different $\ell_{0}$.
@@ 5 Analysis
In this section we investigate the properties of our architecture and of learned features, and compare it to TopK architecture. We show that pre-latents interactions are closely resemble the logic AND gate.
5.1 Toy Model of Superposition
Figure 5: We generate data with covariance matrix that consist of blocks with different sizes on diagonal and off diagonal (left panel). We then examine the decoder-weight covariance $W_{dec}\cdot W_{dec}^{T}$ to assess feature-embedding correlations. Panels (middle panels) show that a TopK SAE recovers these correlation structures only weakly, even after optimal atom matching. In contrast, KronSAE (right panel) more accurately reveals the original block patterns.
To evaluate how well different sparse autoencoder architectures recover known correlation patterns, we construct a controlled experiment using a synthetic, block-structured covariance model. Input vectors $\mathbf{x}_{\mathrm{sparse}}\in\mathbb{R}^{256}$ sampled under a heavy-tailed Bernoulli distribution with probability $p=0.875$ that value will be zero. We perform a Cholesky decomposition $S=LL^{\top}$ on the covariance matrix $S$ and set $\bar{\mathbf{x}}_{\mathrm{sparse}}=L\,\mathbf{x}_{\mathrm{sparse}},$ so that $\bar{\mathbf{x}}_{\mathrm{sparse}}$ exhibits the desired structure.
We train autoencoder (AE) to reconstruct $\mathbf{x}_{\mathrm{sparse}}$. For this we use small toy model following the (Elhage et al., 2022):
$\mathbf{\hat{x}}=\operatorname{ReLU}(W^{T}W\cdot\bar{\mathbf{x}}_{\mathrm{sparse}}+\mathbf{b}).$ (4)
We collect hidden states ($W\cdot\bar{\mathbf{x}}_{\mathrm{sparse}}$) from AE and then train TopK SAE and our proposed KronSAE with $|F|=256$ and $\text{topk}=4$ to reconstruct it. After training, we extract the decoder weight matrices $W_{\mathrm{dec}}$ from each model and compute the covariance $C_{\mathrm{dec}}=W_{\mathrm{dec}}W_{\mathrm{dec}}^{\top}$. To compare TopK SAE embeddings to the AE reference, we match atoms by minimal Euclidean distance, ensuring fair alignment before analysis. Result is shown in Figure 5.
[…]
@@ 5.2 Correlations in SAEs Trained on Language
To examine the correlation structure of features learned in our SAE, we have calculated the correlations on 5k texts from the training dataset. For each feature we calculate the mean correlation with features within its group (head) and with all other features, and compare the randomly initialized KronSAE with $m=4,n=4$ with the trained one. It turns out, as shown in Figure 6, that correlations are indeed significantly higher within a single head and higher than for random SAE, which suggest that our choice to impose the correlated structure in SAE latents works as intended.
@@ 5.3 Analysis of Learned Features
We have compared our KronSAE and TopK architecture in terms of interpretability and feature properties, and we have analyzed the properties of groups in KronSAE. For this, we choose the 14th layer of Qwen2.5-1.5B and a dictionary size of 32k features, of which the first 3072 were selected. KronSAE was chosen with $m=4,n=4$. We run for 24M tokens total to collect data. Our interpretation pipeline follows the common methodology: LLM interprets the activation patterns (Bills et al., 2023) and we evaluate obtained interpretations using the detection score and the fuzzing score Paulo et al. (2024).
For each selected feature, among the standard mean activation value and frequency, we calculate two additional metrics. Low values of token entropy suggest that feature activates more frequently on small number of tokens, thus it is token-specific; high value of multitoken ratio indicates that feature tends to activate multiple times in a single sentence. We have observed that both these metrics have non-negligible negative correlation with the final interpretability scores.
For more details on the data collection and interpretation pipeline, see Appendix D. For additional analysis of properties of learned features, see Appendix E.
@@
Figure 7: Distribution of properties for TopK SAE and KronSAE ($m=4,n=4$) with 32k dictionary size trained on Qwen2.5-1.5B. Pre and Post suffixes denote pre- and post-Cartesian product latents. Our SAE achieves better interpretability scores by learning specialized feature groups, indicated by lower activation frequency and lower variance in activated tokens.
@@
Each post-latent vector has a vector representation in the residual stream represented by the corresponding column in $W_{\text{dec}}$, which is the approximation of overcomplete basis vectors we search for when training SAEs. We had not observed any notable differences in feature geometry between TopK and our SAEs, except that our architectural design leads to clustering so that post-latents produced by same head, base or a extension elements are grouped in a tight cluster, and the geometry is dependent on hyperparameters $h,m,n$ we choose, which is expected and may be useful for further applications such as steering. See more details in Appendix E.
@@ 6 Conclusion and Future Work
We introduce KronSAE, a sparse autoencoder architecture design that directly tackles the long-standing encoder inefficiency bottleneck by combining head-wise Kronecker factorisation with a $\operatorname{mAND}$ gating. This significantly lowers number of parameters compared to standard TopK SAEs, while improving reconstruction fidelity and yielding more interpretable features by utilizing their correlations. Our analysis links these gains to the complementary effects of compositional latent structure and logical AND-style interactions, offering a new lens on how sparsity and factorisation can synergise in representation learning.
@@
Despite its benefits, KronSAE has several limitations. Its gains depend on careful tuning of $(m,n,h)$ and the mAND activation—misconfigured settings can negate both efficiency and quality improvements. Our evaluation is limited to mid-sized transformer models and a single web corpus, so its applicability to larger models, other domains, or languages remains untested.
@@
We identify three directions for extending this work: (i) Transcoding. Treat transcoders (Dunefsky et al., 2024) as implicit routers of information and investigate alternative logical gating functions (e.g. XOR or composite gates) to improve interpretability and circuit analysis. (ii) Crosscoding. Generalize KronSAE to a crosscoder setting (Lindsey et al., 2024) uncover interpretable, cross-level compositionality via logic operations. (iii) Dynamic Composition. Explore learnable tuning of both the number of attention heads and their dimensionality, enabling fine-grained decomposition into groups of correlated features at varying scales.
Appendix A More Details About Experiments
A.1 Experimental setup.
@@ Appendix A More Details About Experiments
All SAEs are optimized using AdamW with an initial learning rate of $8\times 10^{-4}$, a cosine learning-rate schedule with a minimum LR of $1\times 10^{-6}$, and a linear warmup for the first 10% of total training steps. We use a global batch size of 8,192. We sweep over dictionary (latent) sizes of $2^{15}$, $2^{16}$, and $2^{17}$ features. For our KronSAE variant, we further sweep the number of heads $h$ and the per-head dimensions $m$ and $n$ such that $h\cdot m\cdot n$ equals the desired dictionary size. Regularization weights and auxiliary loss coefficients are kept constant throughout the runs to isolate the impact of architectural choices.
@@
As observed by Gao et al. [2025], initializing the decoder as the transpose of the encoder ($W_{\mathrm{dec}}=W_{\mathrm{enc}}^{\top}$) provides a strong metric improvement. We adopt this strategy within KronSAE by partitioning $W_{\mathrm{enc}}$ into $h$ head-wise blocks of shapes $m\times d$ and $n\times d$, denoted $\{P_{i},Q_{i}\}_{i=1}^{h}$. For each head $k$, we define its decoded rows via a simple additive composition:
@@
A.2 FLOPs Calculation.
@@
We calculate flops for most effective variant of TopK where we perform vector matrix multipication only for nonzero activations.
A.3 Pythia model.
For Pythia-1.4B we train all SAEs on the 12th transformer layer. Figure 8 shows that, under larger equivalent token budgets (500M and 1000M), KronSAE achieves a wider performance margin over the TopK SAE than we observed with Qwen-2.5.
Figure 8: Isoflops comparison of KronSAE and TopK SAE on Pythia-1.4B across three token budgets.

### v3

@@ 4.2 Absorption
The notorious challenge in SAE interpretability is feature absorption, where one learned feature becomes a strict subset of another and consequently fails to activate on instances that satisfy the broader concept but not its superset representation (e.g. a “starts with L” feature is entirely subsumed by a “Lion” feature) (chanin2024absorption).
Figure 4 reports three absorption metrics measured via SAEBench (karvonen2025saebench) across sparsity levels $\ell_{0}\in\{16,32,64,128\}$: (1) the mean absorption fraction, measuring the proportion of features that are partially absorbed; (2) the mean full‐absorption score, quantifying complete subsumption events; and (3) the mean number of feature splits, indicating how often a single conceptual feature fragments into multiple activations. We use with dictionary size $F=2^{16}$ and compare TopK SAE, Matryoshka SAE (bussmann2025learning), and their KronSAE version, since they impose different hierarchical priors (see Appendix C for a discussion of how Matryshka SAE and TopK SAE structure differs). Across all $\ell_{0}$, KronSAE variants consistently reduce first two scores relative to the TopK SAE baseline, while maintaining a similar rate of feature splits.
Figure 4: Feature absorption metrics on Qwen-2.5 1.5B and Gemma-2 2B. KronSAE configurations (various $m,n$) exhibit lower mean absorption fractions and full‐absorption scores across different $\ell_{0}$ and selected baselines.
We attribute KronSAE’s improved disentanglement to two complementary design choices:
@@
    AND-like behaviour. By ensuring that post-latent emits only when its more general pre-latent parents are active, we prevent more specific post-latents from entirely subsuming broadly polysemantic one. See additional description of this mechanism in Appendix C.
@@
    Head‐wise Cartesian decomposition. Dividing the latent space into $h$ independent subspaces (each with its own $m\times n$ grid of primitive interactions) ensures that specialized concepts (such as “elephant”) are confined to a single head and cannot fully absorb more general concepts (such as “starts with E”) in another.
Together, these mechanisms produce more monosemantic features, as we verify in the section 5.3, simplifying downstream applications. See results with dictionary size $F=2^{15}$ in Appendix A.4. We also validate the result on Pythia models and observe the same picture, see the Appendix A.5.
@@ 5 Analysis
In this section we examine the properties of the latents in KronSAE compared to TopK architecture.
The design of our arhictecture was also inspired by the observation that many features within a single SAE correlate with each other. KronSAE introduces structural bias that forces the post-latents to co-occur with their pre-latents (see also Appendix C). In this section we analyse if it is helpful.
By examining the toy examples with manufactured correlations in data, we show that KronSAE captures these correlations better than TopK. Then we show that KronSAE trained on language indeed moves correlated features within a single head, indicated by higher correlation within head. After that, we show that KronSAE pre-latents interactions are closely resemble the logic AND gate, and its post-latents are notably more interpretable than TopK latents.
5.1 Toy Model of Correlation
To evaluate how well different sparse autoencoder architectures recover underlying correlation patterns, we construct a controlled experiment using a synthetic, block‐structured covariance model. Input vectors $\mathbf{x}\in\mathbb{R}^{F}$ sampled from a normal distribution (with $\mu=0,\sigma=1$). We then perform a Cholesky decomposition $S=LL^{\top}$ on the covariance matrix $S$ and set $\bar{\mathbf{x}}_{\mathrm{sparse}}=L\,\operatorname{TopK}(\operatorname{ReLU}(\mathbf{x}))$, so that $\bar{\mathbf{x}}_{\mathrm{sparse}}$ exhibits the desired structure.
We train autoencoder (AE) to reconstruct $\bar{\mathbf{x}}_{\mathrm{sparse}}$ following the (elhage2022toymodels):
$\mathbf{\hat{x}}=\operatorname{ReLU}(W^{\top}W\cdot\bar{\mathbf{x}}_{\mathrm{sparse}}+\mathbf{b}).$ (5)
We collect hidden states of dimension $d=64$ ($W\cdot\bar{\mathbf{x}}_{\mathrm{sparse}}$) from AE and then train TopK SAE and our proposed KronSAE with $F=256$ and $\text{topk}=8$ to reconstruct it. After training, we extract the decoder weight matrices $W_{\mathrm{dec}}$ from each SAE, match its latents with the autoencoder latents by solving the quadratic assignment problem (Vogelstein2015FastAQ) (see Appendix B for motivation and details), and compute the covariance $C_{\mathrm{dec}}=W_{\mathrm{dec}}W_{\mathrm{dec}}^{\top}$. Result is shown in Figure 5.
Figure 5: We generate data with covariance matrix that consist of blocks with different sizes on diagonal and off diagonal (left panel). We then examine the decoder‐weight covariance $W_{\mathrm{dec}}\cdot W_{\mathrm{dec}}^{\top}$ to assess feature‐embedding correlations and compute the RV score to quantify the similarity between learned and ground truth covariance matrices. Second panel show feature embeddings for trained autoencoder $W_{\mathrm{enc}}\cdot W_{\mathrm{enc}}^{T}$. Third panel demonstrates that a TopK SAE recovers these correlation structures weakly, as indicated by relatively low RV coefficient $(0.157)$ even after optimal atom matching. In contrast, KronSAE (right panel) more accurately reveals the original block patterns.
To quantify how closely each model’s feature correlations mirror the original structure of $S$, we employ the RV coefficient, defined as $RV(S,C)=\mathrm{trace}(SC)/\sqrt{\mathrm{trace}(S^{2})\,\mathrm{trace}(C^{2})}$. In our experiments, KronSAE consistently achieves notably higher RV than TopK SAE, indicating that our compositional encoder more faithfully reconstructs the original feature relation. See also additional experiments in appendix˜B, where we provide further empirical intuition for how KronSAE identifies correlation structure that more closely align with those present in the data.
@@ 5.2 Correlations in SAEs Trained on Language
To examine the correlation structure of features learned in our SAE, we have calculated the correlations on 5k texts from the training dataset. For each feature we calculate the mean correlation with features within its head and with all other features, and compare the randomly initialized KronSAE with $m=4,n=4$ with the trained one. To isolate the effect of our initialization procedure, we initialize the weights of SAE from the uniform distribution. As shown in Figure 6, correlations are indeed significantly higher within a single head and higher than for random SAE, which suggest that our choice to impose the correlated structure in SAE latents works as intended.
@@ 5.3 Analysis of Learned Features
In this section we compare KronSAE and TopK SAE in terms of interpretability and feature properties, and we analyze the properties of groups in KronSAE. For this, we choose the 14th layer of Qwen2.5-1.5B and a dictionary size of 32k features, of which the first 3072 were selected. KronSAE was chosen with $m=4,n=4$. We run for 24M tokens total to collect data. Our interpretation pipeline follows the common methodology: LLM interprets the activation patterns (bills2023llmcanexplain) and we evaluate obtained interpretations using the detection score and the fuzzing score paulo2024automatically.
For each selected feature, among the standard mean activation value and frequency, we calculate two additional metrics. Low values of token entropy suggest that feature activates more frequently on small number of tokens, thus it is token-specific; high value of multitoken ratio indicates that feature tends to activate multiple times in a single sentence. We have observed that both these metrics have notable negative correlation with the final interpretability scores and therefore they provide useful signal to assess the potential score without calculating it.
For more details on the data collection and interpretation pipeline, see Appendix D. For additional analysis of properties of learned features , additional comparison with baselines and discussion about tradeoff between reconstruction performance and interpretability, see Appendix E.
@@
Figure 7: Distribution of properties for TopK SAE and KronSAE ($m=4,n=4$) with 32k dictionary size trained on Qwen2.5-1.5B. Pre and Post suffixes denote pre- and post- latents, and y-axis indicate density. Token entropy shows the entropy of the distribution of tokens on which feature has activated, and the multitoken ratio measures how often does feature activate in a single sequence.
Our SAE achieves better interpretability scores by learning specialized feature groups, indicated by lower activation frequency and lower variance in activated tokens.
@@
Each post-latent vector has a vector representation in the residual stream represented by the corresponding column in $W_{\text{dec}}$, which is the approximation of overcomplete basis vectors we search for when training SAEs. Our architectural design leads to clustering of feature embeddings so that post-latents produced by same head, base or a extension elements are grouped in a tight cluster, and the geometry is dependent on hyperparameters $h,m,n$ we choose, which is expected and may be useful for further applications such as steering. See more details in Appendix E.
@@ 6 Conclusion and Future Work
We introduce KronSAE, a sparse autoencoder architecture design that combines head-wise Kronecker factorization of latent space with a approximation of logical AND via $\operatorname{mAND}$ nonlinearity. Our approach allows to efficiently train interpretable and compositional SAE, especially in settings with limited compute budget or training data, while maintaining reconstruction fidelity and yielding more interpretable features by utilizing their correlations. Our analysis links these gains to the complementary effects of compositional latent structure and logical AND-style interactions, offering a new lens on how sparsity and factorization can synergise in representation learning.
@@
KronSAE introduces tradeoff between interpretability, efficiency and reconstruction performance, and due to reduced number of trainable parameters it is expected to lag behind TopK SAE at large budgets. Our evaluation is limited to mid-sized transformer models and moderate dictionary sizes; however, the main bottleneck there might be not the SAE itself, but the infrastracture required to handle these setups and the model inference.
@@
We identify three directions for extending this work: (i) Transcoding. Treat transcoders (dunefsky2024transcoders) as implicit routers of information and investigate alternative logical gating functions (e.g. XOR or composite gates) to improve interpretability and circuit analysis. (ii) Crosscoding. Generalize KronSAE to a crosscoder setting (lindsey2024crosscoders) uncover interpretable, cross-level compositionality via logic operations. (iii) Dynamic Composition. Explore learnable tuning of both the number of attention heads and their dimensionality, enabling fine-grained decomposition into groups of correlated features at varying scales.
Ethics Statement
While interpretability research has dual-use potential, our method operates within the ethical boundaries of the underlying models and aims to advance responsible AI development through better model understanding. We analyze activations from publicly available language models (Qwen-2.5-1.5B, Pythia-1.4B, and Gemma-2-2B) gathered on FineWeb-Edu datasets, which excludes the unreported harmful content. We declare no conflicts of interest and maintain transparency about limitations, including potential artifacts from LLM-based interpretation as noted in Appendices D.3 and I.
Reproducibility Statement
We have taken several measures to ensure the reproducibility of our results. We use publicly available models (Qwen, Gemma, Pythia families) and training dataset (FineWeb-Edu) in our experiments. Section 4 and Appendix A provide detailed description of SAE training procedure and hyperparameter configuration. Our complete implementation is available in the supplementary materials, containing the training code, interpretation pipeline and analysis of the results. Appendix H includes simplified implementation of KronSAE that might be easily integrated into existing training codebases, while Appendix D details the interpretability analysis methodology with precise evaluation protocols.
Appendix A Additional Details and Results
A.1 Experimental setup
@@
All SAEs are optimized using AdamW with an initial learning rate of $8\times 10^{-4}$, a cosine learning‐rate schedule with a minimum LR of $1\times 10^{-6}$, and a linear warmup for the first 10% of total training steps, auxiliary loss penalty equal to $0.03125$. We use a global batch size of 8,192. We sweep over dictionary (latent) sizes of $F=2^{15}$, $F=2^{16}$, and $F=2^{17}$ features. For our KronSAE variant, we further sweep the number of heads $h$ and the per‐head dimensions $m$ and $n$ such that $h\cdot m\cdot n$ equals the desired dictionary size. Regularization weights and auxiliary loss coefficients are kept constant throughout the runs to isolate the impact of architectural choices.
@@
As observed by gao2025scaling, initializing the decoder as the transpose of the encoder ($W_{\mathrm{dec}}=W_{\mathrm{enc}}^{\top}$) provides a strong metric improvement. We adopt this strategy within KronSAE by partitioning $W_{\mathrm{enc}}$ into $h$ head‐wise blocks of shapes $m\times d$ and $n\times d$, denoted $\{P_{i},Q_{i}\}_{i=1}^{h}$. For each head $k$, we define its decoded rows via a simple additive composition:
@@
Matryoshka and Kron-based version.
For Matryoshka SAE of dictionary size $F$ we adopt the following experimental setup. We use most training settings from Training details. For Matryoshka SAE bussmann2025learning we define the parameter of dictionary group $S=[2^{k},2^{k},\cdots\,2^{k+i}\cdots\,2^{n}]$, where $k<n$ and $\sum_{s\in S}s=F$. This is equivalent to nested sub-SAEs with dictionary sizes $\mathcal{M}=\{2^{k},2^{k}+2^{k},2^{k}+2^{k}+2^{k+1},\cdots,F\}$ and in our work we define this SAE as Matryoshka-$\mathbf{2^{k}}$. Training loss for Matryoshka is defined as follows:
[…]
where $\mathbf{f}$ is latent vector and $\mathcal{L}_{aux}$ is auxiliary loss used in gao2025scaling.
For Kron-based Matryoshka SAE we use same setup and same initialization from Initialization section. We choose $m$ and $n$ so that $F\mod mn=0$, ensuring that no heads are shared between groups.
Switch SAE and Kron-based version.
For the Switch SAE architecture proposed by mudide2025efficient, we configure all experiments with 8 experts and use identical initialization from KronSAE. In the Kron-based variant of this SAE, we distribute the heads equally across all experts, resulting in $F/(8mn)$ heads per expert. Hence we have less parameters in every expert and therefore we also reduce FLOPs for encoder.
A.2 FLOPs calculation and efficiency
@@
We calculate FLOPs for most effective variant of TopK where we perform vector matrix multipication only for nonzero activations, while encoder still requires dense matrix multiplication.
We have also measured the wallclock time for forward and backward to examine the scaling. Figure 8 reports scaling for different hidden dimension sizes.
Figure 8: Speed comparision of TopK SAE with KronSAE across different hidden dimensionss. We can see that KronSAE have better scaling properties than SAE with default encoder architecture.
A.3 Scaling on Gemma-2 2B
To examine the method’s generality, we conducted additional reconstruction experiments on Gemma-2 2B under iso-FLOPs settings. As shown in Figure 9, our method achieves comparable or improved reconstruction performance across the evaluated compute budgets.
Figure 9: Performance comparision of KronSAE and TopK SAE under a fixed iso-FLOPs budget. Across sparsity settings, KronSAE typically matches TopK’s reconstruction performance and in some cases slightly outperforms it, while using much fewer trainable parameters.
A.4 Smaller dictionary size
To complement our larger-scale experiments, we further evaluate KronSAE’s performance on a smaller dictionary size $F=2^{15}$ with varying $\ell_{0}=\{16,32,64,128\}$ and equal token budget.
Sparsity.
Following the experimental setup described in Section 4.1, we compare KronSAE against established baselines including TopK SAE, Matryoshka SAE, and Switch SAE . As shown in Figure 10, KronSAE consistently matches or exceeds the reconstruction performance of baseline architectures across all tested sparsity levels, while achieving these results with substantially fewer trainable parameters and FLOPs.
Figure 10: Performance comparision of KronSAE and TopK SAE under a fixed token budget. Across sparsity settings, KronSAE typically matches TopK’s reconstruction performance and in some cases slightly outperforms it, while using much fewer trainable parameters and smaller FLOPs budget.
Feature absorption.
We extend our feature absorption analysis from Section 4.2 to the smaller dictionary size, employing two distinct grouping schedules: $S_{1}=[2048,2048,4096,8192,16384]$ and $S_{2}=[1024,1024,2048,4096,8192,16384]$. As shown in Figure 11 KronSAE modification demonstrates better performance, reducing feature absorption metrics.
Figure 11: Feature absorption scores comparison of KronSAE vs baseline methods (TopK SAE and Matryoshka SAE) and a Kron-enhanced Matryoshka SAE variant.
A.5 Pythia Suite
For Pythia-1.4B we train all SAEs on the 12th transformer layer with a budget of 125M tokens. As reported in Table 3, KronSAE achieves performance comparable to TopK SAE with increased number of heads.
Dictionary SAE Mean Max
32k TopK 0.793
KronSAE 0.783 0.793
65k TopK 0.802
KronSAE 0.795 0.805
131k TopK 0.801
KronSAE 0.800 0.810
Table 3: Performance of Pythia-1.4B at 125M budget. At larger dictionary size and fixed training budget KronSAE outperforms TopK SAE.
We conducted additional experiments with smaller Pythia models and trained KronSAE at the middle layers with 512 heads. Table 4 reports results for 125M budget on 65k and 262k dictionary sizes.
Model 70M (d=512) 160M (d=768) 410M (d=1024)
Dictionary 65k 256k 65k 256k 65k 256k
$m=1$ 0.893 0.892 0.856 0.855 0.834 0.835
$m=2$ 0.896 0.897 0.859 0.859 0.832 0.841
$m=4$ 0.894 0.899 0.857 0.859 0.828 0.839
$m=8$ 0.896 0.899 0.856 0.862 0.827 0.846
TopK 0.905 0.903 0.870 0.867 0.843 0.847
Table 4: Performance of SAEs on 70M, 160M and 410M Pythias with varying hidden dimensionality.
In section 4.2 we have analysed whether KronSAE achieves lower absorption score and have answered affirmatively. We also compare the results for Pythia 1.4B model and validate the improvements, as reported in the Table 5.
Model SAE $\ell_{0}=16$ $\ell_{0}=32$ $\ell_{0}=64$ $\ell_{0}=128$ $\ell_{0}=256$
Pythia TopK 0.445 0.233 0.058 0.006 0.004
KronSAE 0.244 0.129 0.033 0.007 0.003
Table 5: Absorption score calculated for Pythia 1.4B model. KronSAE shows lower score due to structured latent space and hierarchy between pre-latents and post-latents.
These results confirm that our compositional architectures improves the absorption score and feature consistency across various models from different families.
A.6 Comparison with JumpReLU
We provide experiments to compare KronSAE with an alternative activation mechanism, JumpReLU, and report explained variance under three sparsity levels in Table 6.
Model Variant $\ell_{0}=32$ $\ell_{0}=50$ $\ell_{0}=64$
TopK 0.809 0.837 0.852
JumpReLU 0.813 0.838 0.844
KronSAE (TopK) 0.814 0.840 0.853
KronSAE (JumpReLU) 0.790 0.817 0.828
Table 6: Performance of SAEs with JumpReLU and TopK activations. Since we have floating sparsity controlled via l0 penalty coefficient, we performed a sweep over various sparsity levels, fitted a parabola to the resulting data as a function of $\ell_{0}$, and evaluated it on those sparsity levels. In contrast, TopK and KronSAE (with TopK) were trained using fixed, predefined sparsity levels.
Replacement of TopK with JumpReLU within KronSAE leads to a degraded performance relative to both JumpReLU SAE and KronSAE with TopK, also with degraded scaling over $\ell_{0}$. This suggests that the architectural advantages of KronSAE interact most effectively with TopK’s behaviour. Whether an alternative activation function can improve on this remains a topic for future work.
A.7 Performance across layers
For this experiment we fix the dictionary size to $F=2^{15}$ and use the same hyperparameters as in the main experiments. As the Figure 12 shows, KronSAE is on-par with TopK at every depth, demonstrating that structured encoder’s reconstruction quality is robust to layer choice.
Figure 12: EV across layers of Qwen-2.5-1.5B, demonstrating that KronSAE matches TopK performance regardless of depth.
A.8 Choice of $m,n,h$
We derive the following guidelines to train KronSAE: one should minimize the $m$ and maximize the $h$ to improve the reconstruction performance, and search for the most expressive configuration from feasible ones. Since $m$ has more impact on EV, one should start from $m=2$ in the search process, since it gives improved computational performance with on-par EV with full TopK training.